The length of a rectangle is 6m greater than its width. If the perimeter must not exceed 80m, calculate the greatest value for its width.
Answer by jorel555(1290)

Let w be the width, l be length:
2w+2l=p, perimeter
2w+2(w+6)=80
4w+12=80
w=17
The greatest width possible is 17m!!!!!!!!!!!!!!!